OpenStudy (anonymous):

The area of the sector formed by the 110 degree central angle is 50 units squared. What is the radius of this circle??

OpenStudy (mathstudent55):

You need the formula for the area of a sector of a circle.

OpenStudy (anonymous):

Solve the following for r: \[\frac{110 \pi r^2}{360}=50 \]\[r=30 \sqrt{\frac{2}{11 \pi }} \]

OpenStudy (anonymous):

Here's The Pic That Goes To My Question

OpenStudy (anonymous):

A complete cicle is 360 degrees and it's area is pi r^2. 110/360 pi r^2 is the area of 110 degrees of the same circle.

OpenStudy (mathstudent55):

@gswag98 This is a problem that needs a formula. You need to either memorize the formula or know where to find it. Then you need to know how to use it. The formula for the area of a sector of a circle is: \(A = \dfrac{\theta}{360^o}\pi r^2\) The Greek letter \(\theta\) (called theta) represents the measure of the central angle that intercepts the arc. In your case, theta = 110 deg. You know the area: A = 50. You are looking for r. The first step is to write the formula, which we have done above. Now replace all the dimensions you were given (theta and A) \(50 = \dfrac{110^o}{360^o}(3.14159) r^2\) \(50 = 0.95993r^2\) \(\dfrac{50}{0.95993} = r^2\) \(r^2 = 52.087\) Now take square root of both sides: \(r = 7.22\)

OpenStudy (anonymous):

is that 7.22 rounded to two decimal plces

OpenStudy (mathstudent55):

Yes.